OpenStudy (anonymous):

I'm studying for a final and I am stumped, I have the answer already but how the heck does one get CosXCosB - SinXSinB from Cos(X + B)?

OpenStudy (anonymous):

\[\cos (\sin^{-1}\left(\begin{matrix}3 \\ 5\end{matrix}\right)+\cos^{-1} \left(\begin{matrix}5 \\ 13\end{matrix}\right))\]

OpenStudy (primeralph):

Well, it's an identity.

OpenStudy (anonymous):

yes

OpenStudy (primeralph):

What does your next statement have to do with your original post?

OpenStudy (anonymous):

that is the question it says find the exact value of that equation

OpenStudy (primeralph):

Is that 3/5 and 5/13?

OpenStudy (anonymous):

yes

OpenStudy (primeralph):

Okay. Now just find a relationship between the double angle formula and your problem. I've found one already. Try.

OpenStudy (anonymous):

i have a paper of all the identities i feel like a complete jackass it is listed there, i have to open my eyes lol the identity states \[\cos (U \pm V) = cosUcosV \mp sinUsinV\]

OpenStudy (primeralph):

Hold, on. I'm helping someone else.

OpenStudy (anonymous):

thanks pal

OpenStudy (primeralph):

Okay, let's start from the top.

OpenStudy (anonymous):

its done man it was a sum and difference identity

OpenStudy (anonymous):

it was was on my sheet of ids

OpenStudy (primeralph):

So you don't need help anymore?

OpenStudy (anonymous):

no you helped me by saying it was an identity I looked at the sheet and there it was lol I wasn't looking hard enough so thank you

OpenStudy (primeralph):

Well, you're smart. Good luck.

OpenStudy (anonymous):

thanks